OpenStudy (anonymous):

tommy can pain a fence in 8 hours. If huck helps, they paint the fence in 1 hour. How fast can huck paint alone?

OpenStudy (anonymous):

1/1 = 1/8 + 1/H -> 1/H = 1 - 1/8 = 7/8 => H = 8/7 = 1 hr 8 min